# Minimal Degrees, Volume Growth, and Curvature Decay on Complete Kähler Manifolds

### v1

@@ Abstract.
We consider noncompact complete Kähler manifolds with nonnegative bisectional curvature. Our main results are: 1. Precise estimates among refined minimal degree of polynomial growth holomorphic functions and holomorphic volume forms, $\operatorname{AVR}$ (asymptotic volume ratio) and $\operatorname{ASCD}$ (average of scalar curvature decay) are established. 2. The Lyapunov asymptotic behavior of the Kähler-Ricci flow can be described in terms of polynomial growth holomorphic functions. This provides a unifying perspective that bridges the two distinct proofs of Yau’s uniformization conjecture. These resolve two conjectures made by Yang.
@@ 1. Introduction
Conjecture 1.1 (Yau’s Uniformization conjecture, [29]).
@@
In the maximal volume growth case, Liu’s breakthrough [16] confirmed the conjecture by combining Gromov-Hausdorff convergence techniques with the three-circle theorem developed in [13]. An alternative proof, based on the Kähler-Ricci flow, was later provided by Lee-Tam [8], building on results in Chau-Tam [3].
These advances led to a deeper understanding of the structure of such manifolds. Notably, through the works of [20, 21, 24, 11, 12, 17], the following conjecture of Ni has been established as a theorem:
Theorem 1.2 (Corollary 3.2 in [20], Theorem 1.2 in [24], Theorem 2 in [12], Theorem 1.4 in [11], Corollary 2.16 in [17]).
@@
$\operatorname{AVR}(M,g)=\lim_{r\to\infty}\frac{\operatorname{Vol}\left(B(p,r)\right)}{\omega_{2n}r^{2n}}=\nu>0.$
@@
$\operatorname{ASCD}(M,g)=\limsup\limits_{r\to\infty}r^{2}\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,r)}S$
@@
(2) A very recent result of Liu [18] implies that for any $(M,g)$ with nonnegative bisectional curvature, the $\limsup$” in the definition of $\operatorname{ASCD}(M,g)$ above can be replaced by $\lim$”, see also [22].
In his thesis [27], Yang defined:
Definition 1.4.
@@
where $\mathcal{K}_{M}$ is the canonical line bundle on $M$, $\mathcal{O}_{P}(M)$ is the space of all polynomial growth holomorphic functions on $M$ and $H^{0}_{P}(M,\mathcal{K}_{M})$ is the space of all polynomial growth holomprhic $n$-forms on $M$. Set $d_{\operatorname{min}}=+\infty$ or $D_{\operatorname{min}}=+\infty$ if $M$ does not admit any nonconstant holomorphic functions of polynomial growth or $\mathcal{K}_{M}$ admits no nonzero holomorphic section of polynomial growth.
Now assume $(M^{n},\omega)$ satisfies the assumption in Theorem 1.2 and of maximal volume growth. Since the Kodaira dimension $K(M^{n})=n$, $\mathcal{O}_{P}(M)$ is “holomorphically regular” in the sense that we can always find “local coordinate by global polynomial growth functions”. The refined minimal degree can be also defined in [27] by
@@
where the infimum is taken among any $n$-tuple of global holomorphic functions that gives local coordinates at $p$ with the corresponding $\limsup_{x\rightarrow\infty}\frac{\log|f_{i}(x)|}{\log d(x,p)}$ arranged in a non-decreasing order for $1\leq i\leq n$, we denote this set by $\mathcal{O}_{P,p}(M)$. In other words, for any $1\leq k\leq n$ we have $d_{\operatorname{min}}^{(k)}=\inf_{f_{k}}\limsup_{x\to\infty}\frac{\log|f_{k}(x)|}{\log d(x,p)}$, where the infimum is taken among all possible $f_{k}$ that appears in the $k$-th component of some sequence in $\mathcal{O}_{P,p}(M)$. Note that apriori it’s unclear if $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ can be obtained by an n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$. But obviously $d_{\operatorname{min}}=d_{\operatorname{min}}^{(1)}$.
Then Yang proposed the following conjecture in [27] on the relation between the above quantities which can be understood as the quantitative version of Theorem 1.2.
Conjecture 1.5 (Conjecture 2.5.6, Conjecture 2.5.8 in [27]).
@@
(1) $\overrightarrow{d_{\operatorname{min}}}(p)$ can be realized by a n-tuple holomorphic functions in $\mathcal{O}_{P,p}(M)$ and is independent of the choice of $p$.
@@
In the case of nonmaximal volume growth, (1),(2) are true because $d_{\min}=+\infty$ from Theorem 1.2, and $\operatorname{ASCD}=+\infty$.
Assume $M$ is of maximal volume growth, in [27], Yang verified the conjecture is true for $U(n)$-invariant Kähler metrics on $\mathbb{C}^{n}$. Liu proved (1) implicitly in [16]. In fact, he showed that any n-tuple of polynomial growth functions which are algebraically independent and of minimal degrees can form a global coordinate. Then recently in [18], Liu proved (2) and derived the explicit formula $\operatorname{ASCD}(M,g)=4n(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$ by passing the geometric quantities to the tangent cones and exploiting the metric Kähler cone structure of the limit space. In [28], Yang proved that (3) holds for gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature by using the Poincaré coordinate introduced in [1], see Theorem 1.8.
Therefore, the general proof of (3) remained the final step toward resolving Conjecture 1.5 in full generality.
In this paper, we prove that (3) of Conjecture 1.5 is true, thereby completely resolving the conjecture. Precisely, we proved
@@
$\frac{\partial g(t)}{\partial t}=-\operatorname{Ric}(g(t)),\quad g(0)=g.$ (1.1)
When the initial metric $(M,g)$ has bounded curvature and Euclidean volume growth, the long-time behavior of (1.1) has been described very clearly. Let us summarize these results in the following:
Theorem 1.6 (Theorem 1.2, Proposition 3.2 in [3], Theorem 6.1 in [4]).
Suppose that $(M^{n},J,g)$ is a complete noncompact Kähler manifold, where $g$ has bounded nonnegative bisectional curvature and Euclidean volume growth with $\operatorname{AVR}(M,g)=\nu>0$. Then the following conclusions hold:
(1) The solution to the Kähler-Ricci flow with the initial metric $g$ exists for all $t\in[0,+\infty)$ and it has nonnegative bisectional curvature for any $t\geq 0$. Moreover, there exists a constant $C(n,\nu)$ such that $\|\operatorname{Rm}(x,t)\|\leq\frac{C(n,\nu)}{t}$ for any $x\in M$ and $t\in(0,+\infty)$.
(2) For any point $x\in M$, let $\{\lambda_{1}(x,t),\cdots,\lambda_{n}(x,t)\}$ denote the eigenvalues of Ricci curvature $\operatorname{Ric}(x,t)$ with respect to $g(t)$ in the nondecreasing order. Then $t\lambda_{i}(x,t)$ is nondecreasing on $t>0$, hence $\mu_{i}(x):=\lim_{t\rightarrow+\infty}t\lambda_{i}(x,t)$ exists.
@@
$\sum\limits_{i=1}^{l}-\mu_{i}(x)\operatorname{dim}_{\mathbb{C}}V_{i}=\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}.$ (1.2)
@@
Indeed, in (3) in Theorem 1.6, $\mu_{i}$ play the role of Lyapunov exponents in a dynamical-system interpretation on asymptotics of (1.1), it basically says that the Ricci curvature can be “simultaneously diagonalized” near $t=+\infty$ in some sense. See Theorem 4.1 in [3] for a precise statement.
Then Yang proposed the following conjecture in [27]:
Conjecture 1.7 (Conjecture 2.5.16 in [27]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with bounded nonnegative curvature and Euclidean volume growth. Let $g(t)$, $t\in[0,+\infty)$, be the complete solution Kähler-Ricci flow with initial metric $g$, then
@@
In [28], Yang proved (1) using Shi’s curvature estimates. He further proved (2) for expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Precisely,
Theorem 1.8 (Theorem 1.4 in [28]).
Let $(N^{n},J,O,g,f)$ be a complete noncompact gradient Kähler-Ricci soliton with nonnegative Ricci curvature, normalized so that $f=R+|\nabla^{g}f|^{2}$. Let $\mu_{1}\leq\cdots\leq\mu_{n}$ be the eigenvalues of Ricci curvature at $O$. Then
$\overrightarrow{d_{\operatorname{min}}}(q)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}\text{ for all }q\in N;$ (1.4)
@@
$D_{\operatorname{min}}=\sum\limits_{i=1}^{n}\mu_{i}.$ (1.5)
If $(M,g)$ has unbounded curvature, in [8], Lee-Tam established the following long-time existence result for the Kähler-Ricci flow:
Theorem 1.9 (Corollary 1 in [21], Theorem 1.5 in [8]).
Suppose $(M^{n},g)$ is a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, $\operatorname{AVR}(M,g)=\nu>0$. Then there exists $C=C(n,\nu)>0$ depending only on $n,\nu$ such that there exists a smooth Kähler-Ricci flow solution $g(t)$ on $[0,+\infty)$ such that
$\|\operatorname{Rm}\|(g(t))\leq\frac{C(n,\nu)}{t}$
on $M\times(0,+\infty)$.
(2) $g(t)$ has nonnegative bisectional curvature.
(3) $\operatorname{AVR}(g(t))=\operatorname{AVR}(g(0))=\nu$ on $M\times(0,+\infty)$.
We emphasize that the uniqueness of such Kähler-Ricci flow solution remains unknown in this setting. Consequently, though these $\mu_{i}$’s can be defined for $M$ with unbounded curvature, a priori, we cannot guarantee that they are independent of the particular flow solution g(t).
In this paper, we prove the Conjecture 1.7 holds in general.
@@
Conjecture 1.7 holds.
Since Conjecture 1.7 is true, then $\mu_{i}$’s are indeed independent of $g(t)$ and thus well-defined for manifolds with unbounded curvature. Therefore we can remove the boundedness condition in Conjecture 1.7.
Corollary 1.10.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative curvature and Euclidean volume growth. Let $g(t)$ $(t\in[0,+\infty)$ be as complete solution Kähler-Ricci flow with initial metric $g$ as in Theorem 1.9, then
(1) $\mu_{i}(p)$ is independent of the choice of $p$ for any $1\leq i\leq n$.
(2) $\mu_{i}=d_{\operatorname{min}}^{(i)}-1$ for any $1\leq i\leq n$.
(3) $\operatorname{ACSD}(M,g(t))$ is invariant along $g(t)$.
Note that by Claim 4.2 in [16], the degrees of functions in $\mathcal{O}_{P}(M)$ also can be “simultaneously diagonalized” along the blowdown sequence, for details see Theorem 2.4 (5) in Section 2. As a by-product of Corollary 1.10, we can describe the Lyapunov regularity in (3) of Theorem 1.6 using the polynomial ring $\mathcal{O}_{P}(M)$. In some sense, it bridges the two distinct proofs of Yau’s uniformization conjecture in the case of maximal volume growth.
Corollary 1.11.
In (3) of Theorem 1.6, the basis of $V_{i}$ can be taken as $\left\{\frac{\partial}{\partial f_{i1}}(x),\cdots,\frac{\partial}{\partial f_{im_{i}}}(x)\right\}$. Here $\operatorname{dim}V_{i}=m_{i}$ is the multiplicity of $\mu_{i}$, $\sum\limits_{i=1}^{l}m_{i}=n$, and
@@
Moreover, $\{f_{11},\cdots,f_{1m_{1}},\cdots,f_{l1},\cdots,f_{lm_{l}}\}$ can serve as a biholomorphism from $M$ onto $\mathbb{C}^{n}$.
This paper is organized as follows. Section 2 is some basic preliminary results and some simple conclusions that will be used later. In section 3, we will prove Theorem B by noting that any tangent cone of $M$ also arises as a tangent cone of corresponding expanding Kähler-Ricci solitons in Theorem 1.6. Also we prove Corollary 1.11. Section 4 contains the proof of Theorem A.
Acknowledgements
I would like to express my sincere gratitude to my advisor, Professor Yihu Yang, for providing constant support. I would like to extend my thanks to Professor Gang Liu for many valuable discussions.

### v2

@@ Abstract.
We consider noncompact complete Kähler manifolds with nonnegative bisectional curvature. Our main results are: 1. Precise relations among refined minimal degree of polynomial growth holomorphic functions and holomorphic volume forms, $\operatorname{AVR}$ (asymptotic volume ratio) and $\operatorname{ASCD}$ (average of scalar curvature decay) are established. 2. The Lyapunov asymptotic behavior of the Kähler-Ricci flow can be described in terms of polynomial growth holomorphic functions. This provides a unifying perspective that bridges the two distinct proofs of Yau’s uniformization conjecture by Liu and Chau-Lee-Tam. These resolve two conjectures made by Yang. 3. the link of any tangent cone is homeomorphic to a sphere, providing a Kähler analogue of a result of Kapovitch in the Riemannian case.
@@ 1. Introduction
Conjecture 1 (Yau’s Uniformization Conjecture, [Yau91]).
@@
In the maximal volume growth case, Liu’s breakthrough [Liu19] confirmed the conjecture by combining Gromov-Hausdorff convergence techniques with the three-circle theorem developed in [Liu16c]. An alternative proof, based on the Kähler-Ricci flow, was later provided by Lee-Tam [LT20], building on results in Chau-Tam [CT06].
These advances lead to a deeper understanding of the structure of such manifolds. Notably, through the works of [Ni04, Ni05, NT13, Liu16a, Liu16b, Liu21], the following conjecture of Ni has been established as a theorem:
Theorem 1 (Corollary 3.2 in [Ni04], Theorem 1.2 in [NT13], Theorem 2 in [Liu16b], Theorem 1.4 in [Liu16a], Corollary 2.16 in [Liu21]).
@@
$\operatorname{AVR}(M,g)=\lim_{r\to\infty}\frac{\operatorname{Vol}\left(B(p,r)\right)}{\omega_{2n}r^{2n}}>0.$
@@
$\operatorname{ASCD}(M,g)=\limsup\limits_{r\to\infty}r^{2}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,r)}S$
@@
(2) A very recent result of Liu [Liu24] shows that for any $(M,g)$ with nonnegative bisectional curvature, the $\limsup$” in the definition of $\operatorname{ASCD}(M,g)$ above can be replaced by $\lim$”, see also [Ni12].
In his thesis [Yang13], Yang defined:
Definition 1.
@@
where $\mathcal{K}_{M}$ is the canonical line bundle on $M$, $\mathcal{O}_{P}(M)$ is the space of all polynomial growth holomorphic functions on $M$ and $P(M,\mathcal{K}_{M})$ is the space of all polynomial growth holomorphic $n$-forms on $M$. Set $d_{\operatorname{min}}=+\infty$ or $D_{\operatorname{min}}=+\infty$ if $M$ does not admit any nonconstant holomorphic functions of polynomial growth or $\mathcal{K}_{M}$ admits no nonzero holomorphic sections of polynomial growth.
Now assume $(M^{n},\omega)$ satisfies the assumption in Theorem 1 and is of maximal volume growth. Since the Kodaira dimension $K(M)=n$, $\mathcal{O}_{P}(M)$ is “holomorphically regular” in the sense that we can always find “local coordinate by global polynomial growth holomorphic functions”. The refined minimal degree can be also defined in [Yang13] by
@@
where the infimum is taken among any $n$-tuple of global holomorphic functions that gives local coordinate at $p$ with the corresponding $\limsup\limits_{x\rightarrow\infty}\frac{\log|f_{i}(x)|}{\log d(x,p)}$ arranged in a non-decreasing order for $1\leq i\leq n$, we denote this set by $\mathcal{O}_{P}(M,p)$. In other words, for any $1\leq k\leq n$ we have $d_{\operatorname{min}}^{(k)}=\inf\limits_{f_{k}}\limsup\limits_{x\rightarrow\infty}\frac{\log|f_{k}(x)|}{\log d(x,p)}$, where the infimum is taken among all possible $f_{k}$ that appears in the $k$-th component of some sequence in $\mathcal{O}_{P}(M,p)$. Note that apriori it’s unclear if $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ can be obtained by an n-tuple holomorphic functions in $\mathcal{O}_{P}(M,p)$. But obviously $d_{\operatorname{min}}=d_{\operatorname{min}}^{(1)}$.
Then Yang proposed the following conjecture on the relation between the above quantities which can be understood as the quantitative version of Theorem 1.
Conjecture 2 (Conjecture 2.5.6, Conjecture 2.5.8 in [Yang13]).
@@
(1) If $M$ is of maximal volume growth, then $\overrightarrow{d_{\operatorname{min}}}(p)$ can be realized by a n-tuple holomorphic functions in $\mathcal{O}_{P}(M,p)$ and is independent of the choice of $p$.
@@
In the case of nonmaximal volume growth, (2) is true because $d_{\min}=+\infty$ from Theorem 1. Also we can show that $\operatorname{ASCD}=+\infty$.
Assume $M$ is of maximal volume growth, in [Yang13], Yang verified the conjecture is true for $U(n)$-invariant Kähler metrics on $\mathbb{C}^{n}$. Liu proved (1) implicitly in [Liu19]. In fact, he showed that any n-tuple of polynomial growth functions which are algebraically independent and of minimal degrees can form a global coordinate. Then recently in [Liu24], Liu proved (2) and derived the explicit formula $\operatorname{ASCD}(M,g)=4n(\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n)$ by passing the geometric quantities to the tangent cones and exploiting the metric Kähler cone structure of the limit space. In [Yang22], Yang proved that (3) holds for gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature by analysing the Poincaré coordinate introduced in [Bry08], see Theorem 3 for more details.
Therefore, the general proof of (3) remained the final step toward resolving Conjecture 2 in full generality.
In this paper, we prove that (3) of Conjecture 2 is true, thereby completely resolving the conjecture. Precisely,
@@
(1.1) $\frac{\partial g(t)}{\partial t}=-\operatorname{Ric}(g(t)),\quad g(0)=g.$
When the initial metric $(M,g)$ has maximal volume growth, the long-time behavior of (1.1) has been described very clearly. Let us summarize these results in the following:
Theorem 2 (Corollary 1 in [Ni05], Theorem 1.2, Proposition 3.2 in [CT06], Theorem 6.1 in [CT11], Theorem 1.5 in [LT20], Theorem 1.1 in [Lee25]).
Suppose that $(M^{n},J,g)$ is a complete noncompact Kähler manifold, where $g$ has nonnegative bisectional curvature and maximal volume growth with $\operatorname{AVR}(M,g)=\nu>0$. Then the following conclusions hold:
(1) There exists a unique long-time solution $g(t)$ to the Kähler-Ricci flow on $M\times[0,+\infty)$ with the initial metric $g$ such that,
  1. (i)
    It has nonnegative bisectional curvature for any $t\geq 0$.
  2. (ii)
    There exists a constant $C(n,\nu)$ such that $\|\operatorname{Rm}(x,t)\|\leq\frac{C(n,\nu)}{t}$ for any $x\in M$ and $t\in(0,+\infty)$.
  3. (iii)
    $\operatorname{AVR}(M,g(t))=\operatorname{AVR}(M,g)=\nu$.
(2) For any point $x\in M$, let $\{\lambda_{1}(x,t),\cdots,\lambda_{n}(x,t)\}$ denote the eigenvalues of Ricci curvature $\operatorname{Ric}(x,t)$ with respect to $g(t)$ in the nondecreasing order. Then $t\lambda_{i}(x,t)$ is nondecreasing on $t>0$, hence $\mu_{i}(x):=\lim\limits_{t\rightarrow+\infty}t\lambda_{i}(x,t)$ exists.
@@
(1.2) $\sum\limits_{i=1}^{l}-\mu_{i}(x)\operatorname{dim}_{\mathbb{C}}V_{i}=\lim\limits_{t\to\infty}\log\frac{\det(g_{i\bar{j}}(x,t))}{\det(g_{i\bar{j}}(x,0))}.$
@@
Remark.
(1) If $(M,g)$ has bounded curvature, the above theorem was proved by Chau–Tam in [CT06, CT11] through Li-Yau-Hamilton estimate. In the case of unbounded curvature, the existence of such solution was established by Lee–Tam in [LT20, LT21], in which they constructed a pyramid solution to the Kähler–Ricci flow. Moreover, the uniqueness of complete non-compact Ricci flows under a scaling-invariant curvature bound was proved very recently by Lee [Lee25].
(2) In the context of the Kähler–Ricci flow, the theorem asserts that $(M,J,g(t))$ asymptotically approaches an expanding gradient Kähler–Ricci soliton at the base point $p$. Moreover, conclusions (2) and (3) essentially indicate that the Ricci curvature $\operatorname{Ric}(x,t)$ becomes “simultaneously diagonalizable” as $t\rightarrow+\infty$, in an appropriate asymptotic sense. In fact, $\mu_{i}$’s play the role of Lyapunov exponents in a dynamical-system interpretation on asymptotics of (1.1).
Then Yang proposed the following conjecture in [Yang13]:
Conjecture 3 (Conjecture 2.5.16 in [Yang13]).
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative curvature and maximal volume growth. Let $g(t)$, $t\in[0,+\infty)$, be the unique complete solution Kähler-Ricci flow with initial metric $g$ in Theorem 2, then
@@
In [Yang22], Yang proved (1) using Shi’s curvature estimates on Ricci flow solutions under a scaling-invariant curvature bound. He further proved (2) for expanding Kähler-Ricci solitons with nonnegative Ricci curvature. Precisely,
Theorem 3 (Theorem 1.4 in [Yang22]).
Let $(N^{n},J,O,g,f)$ be a complete noncompact gradient Kähler-Ricci soliton with nonnegative Ricci curvature, normalized so that $f=R+|\nabla^{g}f|^{2}$. Let $0\leq\mu_{1}\leq\cdots\leq\mu_{n}$ be the eigenvalues of Ricci curvature at $O$. Then
(1.4) $\overrightarrow{d_{\operatorname{min}}}(q)=\{\mu_{1}+1,\cdots,\mu_{n}+1\}\text{ for all }q\in N;$
@@
(1.5) $D_{\operatorname{min}}=\sum\limits_{i=1}^{n}\mu_{i}.$
In this paper, we show that the Conjecture 3 holds in general.
@@
Conjecture 3 holds.
Note that by Claim 4.2 in [Liu19], the degrees of polynomial growth functions also can be “simultaneously diagonalized” along the metric blow-down sequence, we will discuss the phenomenon in Theorem 5 (5) in Section 2. As a by-product, we can describe the Lyapunov regularity in (3) of Theorem 2 using the polynomial ring $\mathcal{O}_{P}(M)$. In some sense, it bridges the two distinct proofs of Yau’s uniformization conjecture in the case of maximal volume growth.
Corollary 1.
In (3) of Theorem 2, the basis of $V_{i}$ can be taken as $\left\{\frac{\partial}{\partial f_{i1}}(x),\cdots,\frac{\partial}{\partial f_{im_{i}}}(x)\right\}$. Here $\operatorname{dim}V_{i}=m_{i}$ is the multiplicity of $\mu_{i}$, $\sum\limits_{i=1}^{l}m_{i}=n$, and
@@
Moreover, $\{f_{11},\cdots,f_{1m_{1}};\cdots;f_{l1},\cdots,f_{lm_{l}}\}$ can serve as a biholomorphism from $M$ onto $\mathbb{C}^{n}$.
Now let’s turn our attention to the tangent cone of $(M,g)$ with nonnegative bisectional curvature and maximal volume growth. By Lott [Lott21], any tangent cone of $M$ is a smooth complex manifold which is biholomorphic to $\mathbb{C}^{n}$ with Kähler current $\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2}$, where $r$ is the radial function on the tangent cone. Moreover, it satisfies the condition $\operatorname{BK}\geq 0$”, which is a new pluripotential-theoretic characterization which is equivalent to nonnegative bisectional curvature on Kähler manifolds, and can pass to Gromov-Hausdorff limits.
In this paper, we shall prove the following structure theorem about the tangent cones:
Theorem C.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $(M_{\infty}=C(X),d)$ be a tangent cone at infinity of $M$, which is biholomorphic to $\mathbb{C}^{n}$, with link $X$.
(1) $X$ is homeomorphic to the sphere $\mathbb{S}^{2n-1}$.
(2) If we assume $(M,g)$ has quadratic curvature decay, then there exists a sequence of smooth Kähler cone metric $\omega_{i}=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r_{i}^{2}$, which have the same Reeb field with $M_{\infty}$, such that $(\mathbb{C}^{n},\omega_{i})$ converges to $(M_{\infty},d)$ in the Gromov-Hausdorff sense. In particular, for any fixed point $p\in\mathbb{C}^{n}$, $\operatorname{BK}(\omega_{i},p)\geq-\Psi(\frac{1}{i})$. Here, and below, $\Psi(\frac{1}{i})$ denotes a positive function converging to zero as $i\rightarrow+\infty$.
Remark.
(1) Theorem C (1) can be seen as a Kähler analogue of Theorem 1.3 in Kapovitch [Kap02], which states that the link of any tangent cone of a complete noncompact Riemannian manifold with nonnegative sectional curvature and maximal volume growth has Alexandrov curvature bounded below by $1$ and is homeomorphic to a sphere.
(2) From Theorem 1 we know that if $M$ is of maximal volume growth, then the curvature is of average quadratic decay. So the assumption of (2) doesn’t seem to be that strict. Nevertheless, we still wonder if this condition can be removed. Also, we wonder if we can get a uniform bisectional curvature lower bound, i.e. $\operatorname{BK}(\omega_{i})\geq-\frac{\Psi(\frac{1}{i})}{r_{i}^{2}}$ uniformly.
This paper is organized as follows. Section 2 is some basic preliminary results and some simple conclusions that will be used later. In Section 3, we will prove Theorem B by noting that any tangent cone of $M$ also arises as a tangent cone of corresponding expanding Kähler-Ricci solitons described in Theorem 2. After finishing the proof, we will prove Corollary 1. Section 4 contains the proof of Theorem A. In Section 5 we will prove Theorem C using some knowledge of Sasaki geometry and discuss several further questions.
Acknowledgement: I would like to express my sincere gratitude to my advisor, Yihu Yang, for his constant support and lots of inspiring discussions. I would like to extend my thanks to Gang Liu and Zhiqin Lu for many valuable discussions. I am thankful to Man-Chun Lee for his interest in this work and valuable comments. I am grateful to University of California, Irvine for hospitality during my visit in fall 2025.